1.If a Japanese car costs 500,000 yen, a similar American car costs $10,000, and a dollar can buy 100 yen, what are the nominal and real interest rates?


1
Expert's answer
2021-08-16T08:47:54-0400


One dollar is buying 100 yen. Thus, nominal exchange rate is $1 that equal to 100 yen.

Real exchange rate can be calculated as follows.

Real interest rate"=\frac{car \space price\space in\space america}{car\space price \space in\space japan}\times exchange \space rate\space yen\space per\space dollar"

"=\frac{10000}{500000}\times100=2"

Real interest rate is1 American car is equal to 2 Japanese car.